Wimpy earns $200 per month and consumes 60 Hardee’s hamburgers each month. He pays $2 for each hamburger. Hardee’s is considering implementing a frequent-eater’s program whereby a monthly charge of $30 allows you to purchase each hamburger for $1.50. If Wimpy chooses to do this will he eat more hamburger or less? Is he better or worse off? Should Wimpy join this frequent-eaters program?
Answer #1

As a regular customer his monthly expenditure on burgers is = 60* 2 = $120

as a frequent eater program, his monthly expenditure is = 30 + ( 1.5*60 ) = 120

if he eat same quantity of burgers per month then there is no change in his monthly expenditure.

but if his increase the number of burgers to eat then he is better off

suppose as frequent eater problem if he eat 3 burgers daily then his expenditures will be

= 30+(1.5*90) = $165

as a regular customer his monthly expenditure for 90 burgers is = 90* 2 = $180

so if he want to consume more then 60 burgers then this offer is good otherwise there is no benefit no lose